# Towards Real-world Video Face Restoration: A New Benchmark

### v1

@@ Abstract
Blind face restoration (BFR) on images has significantly progressed over the last several years, while real-world video face restoration (VFR), which is more challenging for more complex face motions such as moving gaze directions and facial orientations involved, remains unsolved. Typical BFR methods are evaluated on privately synthesized datasets or self-collected real-world low-quality face images, which are limited in their coverage of real-world video frames. In this work, we introduced new real-world datasets named FOS with a taxonomy of “Full, Occluded, and Side" faces from mainly video frames to study the applicability of current methods on videos. Compared with existing test datasets, FOS datasets cover more diverse degradations and involve face samples from more complex scenarios, which helps to revisit current face restoration approaches more comprehensively. Given the established datasets, we benchmarked both the state-of-the-art BFR methods and the video super resolution (VSR) methods to comprehensively study current approaches, identifying their potential and limitations in VFR tasks. In addition, we studied the effectiveness of the commonly used image quality assessment (IQA) metrics and face IQA (FIQA) metrics by leveraging a subjective user study. With extensive experimental results and detailed analysis provided, we gained insights from the successes and failures of both current BFR and VSR methods. These results also challenge current face restoration approaches, which we hope stimulate future advances in VFR research.  [^1] Project page: https://ziyannchen.github.io/projects/VFRxBenchmark/.
@@ 1 Introduction
Figure LABEL:subfig1:previous_samples shows that state-of-the-art BFR methods can yield incredible results in both identity preservation and facial features/texture details regeneration on widely used real image testing datasets (e.g., WebPhoto-Test [48], LFW-Test [18], Wider-Test [61], CelebChild [48]). However, these commonly used test datasets cover limited scenarios. Specifically, WebPhoto-Test and Wider-Test are mainly about human photography, where (1) most faces are assured to have no occlusion, (2) diversity of the gaze and head directions are lacking, and (3) the facial expressions are not diverse enough. Although these missing features may be marginal in the image cases, they are commonly found in video cases. Therefore, when applying video test samples to existing BFR methods, we find their clear limitations in handling complex image faces or video motions. Figure LABEL:subfig1:fos_demo indicates obvious failures of the most recent state-of-the-art BFR method CodeFormer [61] in restoring side/occluded faces or maintaining stability in face video frames.
These observations revealed limitations of the existing image test datasets. Furthermore, real video face test sets remain lacking. Therefore, we introduce a new benchmark to advance future VFR research in this work. Specifically, we construct three test datasets, FOS-V, FOS-real and FOS-syn, with three representative face poses/conditions highlighted as Full, Occluded, and Side. An overview of our FOS datasets is presented in Table 1, while visualized samples are in Figure 2. Compared to the existing test sets, FOS datasets show larger diversity that involves heterogeneous scenes (e.g. interviews, sports, concerts, and old movies) and a wide range of faces with different facial expressions, poses, and ethnicities.
@@
    We introduce a real-world video face dataset, FOS-V, filling a gap in the VFR field where real test sets are missing.
@@
    By benchmarking the state-of-the-art BFR and VSR on FOS, we identify future challenges of FVR task. None of the existing methods can achieve satisfying intra-frame reconstruction performance while maintaining inter-frame stability.
@@
    In addition, we leverage a user study to investigate the effectiveness of IQA and FIQA metrics. The results show that NIQE [37] and BRISQUE [36] are significantly inconsistent with subjective scores, while FIQA [45, 16, 40] metrics yield a tremendously strong correlation with the human eye.
@@ 2 Related Work
Video restoration Although BFR methods have achieved great success in single image restoration, video face restoration draws little attention from researchers. A few attempts have been made in video face super resolution [12, 34, 53], most of which focus on a fusion of inter-frame spatial and temporal information, or aural and visual modalities. Recently, [62] made a new attempt to apply the diffusion model to the VFR task. It is also worth noting that some modest progress has been made in general video super resolution. For example, EDVR [47] and BasicVSR [4] demonstrated their appealing performance in both temporal consistency and single-frame restoration by leveraging temporal alignment, and aggregation.
@@ 4 Experiments
We evaluate 11 representative state-of-the-art BFR methods on FOS-syn and FOS-real: GAN prior-based methods (GFP-GAN [48], GLEAN [3], GPEN [57], PULSE [35]), VQ-based methods (VQFR [14], CodeFormer [61], RestoreFormer [50]), reference-based method (DMDNet [^2] An extension of DFDNet [30]. [32]), and other methods (GCFSR [15], PSFRGAN [5], HiFaceGAN [55]). Furthermore, we evaluate 4 VSR models trained on VFHQ [52](EDVR [47], EDVR-GAN, BasicVSR [4], BasicVSR-GAN) on our real-world FOS-V dataset and the synthesized VFHQ-Test [52]. As a comparison, 6 most recent BFR methods (GFP-GAN [48], GPEN [57], VQFR [14], CodeFormer [61], GCFSR [15], RestoreFormer [50]) are also evaluated on those video datasets in a single-frame processing manner. 333 DMDNet is excluded from the comparison since detecting 68 landmarks often fails under complex conditions.

### v2

@@ Abstract
Blind face restoration (BFR) on images has significantly progressed over the last several years, while real-world video face restoration (VFR) remains unsolved. One current challenge in the VFR task is the absence of an effective evaluation system. The real-world video face test dataset is lacking, and the evaluation metrics require validation and supplementation before applying to video scenarios. In this work, we introduced new real-world datasets named FOS with a taxonomy of “Full, Occluded, and Side" faces derived from mainly video frames to study the applicability of current methods on videos. Compared with existing test datasets, FOS datasets cover real-world video test sets, more diverse degradations, and complex scenarios. Given the established datasets, we benchmarked both the state-of-the-art BFR methods and the video super resolution (VSR) methods to comprehensively study current approaches, identifying their potential and limitations in VFR tasks. In addition, we studied the effectiveness of the commonly used image quality assessment (IQA) metrics and face IQA (FIQA) metrics by leveraging a subjective user study. The extensive experimental results and detailed analysis pose challenges to current face restoration approaches, which we hope stimulate future advances in VFR research. The project page can be found in https://ziyannchen.github.io/projects/VFRxBenchmark/.
@@ 1 Introduction
Despite promising progress in BFR, the video face restoration (VFR) remains unsolved. Although a few methods have been proposed in face super-resolution [62, 12, 34, 53], there is still a huge gap between the performances of BFR and VFR methods. Figure LABEL:subfig1:previous_samples shows that state-of-the-art BFR methods can yield incredible results in both identity preservation and facial features/texture details regeneration on widely used real image testing datasets (e.g., WebPhoto-Test [48], LFW-Test [18], Wider-Test [61], CelebChild [48]). However, these commonly used test datasets cover limited scenarios. Specifically, WebPhoto-Test and Wider-Test are mainly about human photography, where (1) most faces are assured to have no occlusion, (2) diversity of the gaze and head directions are lacking, and (3) the facial expressions are not diverse enough. Although these missing features may be marginal in the image cases, they are commonly found in video cases. Figure LABEL:subfig1:fos_demo indicates obvious failures of the most recent state-of-the-art BFR method CodeFormer [61] in restoring side/occluded faces or maintaining stability in face video frames.
These observations reveal that a complete evaluation system applicable to the VFR task is missing. Real video face test sets remain lacking, and limitations of the existing image test datasets are not fully reflected. Therefore, we introduce a new benchmark in this work, namely FOS, hoping to advance future VFR research. Specifically, we construct three test datasets, FOS-V, FOS-real and FOS-syn, with three representative face poses/conditions highlighted as Full, Occluded, and Side. An overview of our FOS datasets is presented in Table 1, while visualized samples are in Figure 2. Compared to the existing test sets, FOS datasets show larger diversity that involves heterogeneous scenes (e.g. interviews, sports, concerts, and old movies) and a wide range of faces with different facial expressions, poses, and ethnicities.
@@
    We introduce a comprehensive real-world face test dataset, involving a video test subset to fill a gap in the VFR field where real test sets are missing.
@@
    By benchmarking the state-of-the-art BFR and VSR methods, we identify future challenges in FVR field. None of the existing methods can achieve satisfying intra-frame reconstruction performance while maintaining inter-frame stability.
@@
    We leverage a user study to investigate the effectiveness of IQA and FIQA metrics. The results show that NIQE [37] and BRISQUE [36] are significantly inconsistent with subjective scores, while FIQA [45, 16, 40] metrics yield a strong correlation with the human eye.
@@ 2 Related Work
Video restoration Although BFR methods have achieved great success in single image restoration, video face restoration draws little attention from researchers. A few attempts have been made in video face super resolution [12, 34, 53, 62], most of which focus on a fusion of inter-frame spatial and temporal information, or aural and visual modalities. Recently, [62] made a new attempt to apply the diffusion model to the VFR task. It is also worth noting that some modest progress has been made in general video super resolution. For example, EDVR [47] and BasicVSR [4] demonstrated their appealing performance in both temporal consistency and single-frame restoration by leveraging temporal alignment, and aggregation.
@@ 4 Experiments
We evaluate 11 representative state-of-the-art BFR methods on FOS-syn and FOS-real: GAN prior-based methods (GFP-GAN [48], GLEAN [3], GPEN [57], PULSE [35]), VQ-based methods (VQFR [14], CodeFormer [61], RestoreFormer [50]), reference-based method (DMDNet [^1] An extension of DFDNet [30]. [32]), and other methods (GCFSR [15], PSFRGAN [5], HiFaceGAN [55]). Furthermore, we evaluate 4 VSR models trained on VFHQ [52](EDVR [47], EDVR-GAN, BasicVSR [4], BasicVSR-GAN) on our real-world FOS-V dataset and the synthesized VFHQ-Test [52]. As a comparison, 6 most recent BFR methods (GFP-GAN [48], GPEN [57], VQFR [14], CodeFormer [61], GCFSR [15], RestoreFormer [50]) are also evaluated on those video datasets in a single-frame processing manner. 222 DMDNet is excluded from the comparison since detecting 68 landmarks often fails under complex conditions.